# Is More Context Always Better? Examining LLM Reasoning Capability for Time Interval Prediction

### v1

@@ Abstract.
Large Language Models, Temporal Reasoning, Inter-Purchase Interval Prediction
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Recommender systemsccs: Information systems Personalization
@@ 1. Introduction
Predicting the time until a user’s next action is a fundamental yet often overlooked dimension of behavioral modeling on the Web. This interval, whether for replenishing groceries or renewing a subscription, tells us about users’ habits and needs. Accurately forecasting this interval is therefore a key component of creating personalized experiences. Traditional machine learning (ML) and statistical approaches have long been used to estimate these intervals with structured data. However, the recent emergence of LLMs raises a question: can an LLM reason about time intervals as effectively as specialized predictive models?
Though LLMs show strong capability in language understanding, mathematical reasoning and multi-step logical inference, their ability in structured temporal reasoning remains largely unexamined. Most studies focus on what LLMs can infer from text, not on whether they can capture quantitative regularities hidden in behavioral data.
@@
Recent work applies LLMs to temporal forecasting by converting time series into textual sequences. Models such as LLMTime (Gruver et al., 2024) treats numerical time series as digit string tokens, enabling zero-shot forecasting without model modifications. TIME-LLM (Jin et al., 2024) reprograms LLM by learning an alignment layer that translates time series values into its lexical space. Methods like LSTPrompt (Liu et al., 2024) prompt LLMs to reason separately about long and short range patterns. Despite these advances, their benefits over classical models remain uncertain. (Tang et al., 2024) shows that LLMs can handle periodic patterns but struggle with irregular ones, although external context can help. In user modeling, Wei et al. (Du et al., 2025) incorporate irregular time intervals into LLMs for sequential recommendation, yet the task remains focused on what to recommend rather than when the event occurs.
@@ 2.1. Datasets
We evaluate our approach on two real-world e-commerce datasets: proprietary grocery data and the public Instacart dataset (M Yasser H., 2017). In both datasets, user-category pairs with fewer than five purchases are excluded, and intervals longer than 20 days are removed. The proprietary dataset contains 5,780 users across 12 product types (1.03 types/user average) with over 110,000 purchase records. The Instacart dataset includes 2,661 users across 10 categories (1.87 categories/user average) with roughly 194,000 orders. Both datasets span diverse categories (Fresh Produce, Dairy, Packaged Goods, Snacks) capturing variability in purchase periodicity and seasonality for comparing structured and language-based temporal reasoning.
@@ 3.1. Main Results
RQ1: LLMs vs. Traditional ML A primary finding from 1, consistent across both the proprietary and Instacart datasets, is the ML model’s dominant performance. The traditional ML model achieves over 50% improvement in most key metrics compared to the best-performing LLM. For instance, on the proprietary data dataset, the ML model’s MAPE of 29.92% is 92.0% better than the best-performing LLM (Claude-3.5 Medium at 57.45%). For the business-critical ”TA@1” metric, the ML model scores 22.98%, beating the best LLM’s 14.63% (Claude-3.5 High) by 57.1%. This performance gap is likely due to a fundamental ”impedance mismatch” for the LLM. The ML model is a mathematical function approximator fed with structured numerical features. The LLM, in contrast, must first translate qualitative, linguistic descriptions into an internal quantitative representation and then perform a regression. This two-step process introduces significant error. However, the disparity in performance is significantly larger for error metrics than for accuracy metrics, as LLM’s performance on TA@k is competitive. This suggests that while LLMs fail at quantitative precision (pinpointing the exact day), they are better at approximate temporal classification. This aligns with the business goal for replenishment, where knowing a user will buy ”in the next 48 hours” is often sufficient.
@@
Another notable observation is that LLM predictions outperform the ”stat best” median baseline. Across both datasets, LLMs consistently improve upon this baseline on both metrics, demonstrating that they can extract additional signal from contextual descriptions beyond what a simple historical median provides. This indicates that LLMs are not merely parroting the central tendency of the data; rather, they do incorporate contextual cues to produce more informed estimates, especially when the goal is approximate rather than exact-day prediction. These results suggest that LLMs have useful but limited temporal reasoning capabilities.
@@
RQ2: The Impact of Context on LLM Performance Our second research question examines how different levels of context affect LLM performance. Across both datasets, medium-level context—consisting of basic product attributes, simple statistics and concise user history yields consistent improvements in both model metrics and business-oriented accuracy. In contrast, adding high-level narrative and context information often reduces accuracy, sometimes performing no better than the zero-context baseline. This suggests a “context-as-noise” effect: while medium context provides focused, decision-relevant information, high-context prompts introduce semantically rich but temporally irrelevant details that can mislead the model’s internal reasoning. Figure 3 illustrates this dynamic. The green highlights show that medium-context prompts encourage the model to rely on stable quantitative cues such as dominant intervals and simple trends, producing correct or near-correct predictions. The red highlights show that high-context prompts surface additional narrative details that the model overweights, diverting attention from the underlying temporal structure and leading to erroneous predictions.
Finally, when comparing LLM families, Claude-3.5 shows the most consistent advantage, outperforming all rivals on the internal dataset. While on Instacart dataset, both Claude-3.5 and Gemini-2.5 achieve top performance on different sets of metrics. In terms of efficiency, GPT-4o is the fastest and cheapest (around 1.2-1.4s latency, ¿\$0.003 per call), while Claude-3.5 is moderately slower (around 2-4s) and 3× more expensive but still far more efficient than Gemini-2.5, which has the highest latency (around 15-19s) with lower accuracy.
Figure 3. Examples of Claude-3.5-Sonnet prediction and reasoning with different levels of context. Green marks the correct medium level information leveraged, while red marks the additional high context leading to incorrect prediction.
@@
Our study presents a framework for evaluating time-interval prediction, establishes strong baselines, and positions LLMs between statistical and machine-learning models. Our analysis of context levels further reveals a key trade-off: targeted and compact context can aid LLM reasoning, while excessive narrative detail can obscure the underlying temporal signal. These findings highlight both the promise and current limits of LLMs for structured temporal inference and point toward future hybrid models that blend statistical precision with linguistic flexibility.

### v2

@@ Abstract.
Large Language Models, Temporal Reasoning, Sequence Modeling
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792900isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computing methodologies Temporal reasoning
@@ 1. Introduction
Predicting the time until a user’s next action is a fundamental yet often overlooked dimension of sequential behavioral modeling on the Web (Cao et al., 2025; Ma et al., 2025). This interval, whether for replenishing groceries or renewing a subscription, tells us about users’ habits and needs. Accurately forecasting this interval is therefore a key component of creating personalized experiences. Traditional machine learning (ML) and statistical approaches have long been used to estimate these intervals with structured data. However, the recent emergence of LLMs raises a question: can an LLM reason about time intervals as effectively as specialized predictive models?
Though LLMs show strong capability in language understanding, mathematical reasoning and multi-step logical inference  their ability in structured temporal reasoning remains largely unexamined. Most studies focus on what LLMs can infer from text (Zhang et al., ), not on whether they can capture quantitative regularities hidden in behavioral data.
@@
Recent work applies LLMs to temporal forecasting by converting time series into textual sequences. Models such as LLMTime (Gruver et al., 2024) treats numerical time series as digit string tokens, enabling zero-shot forecasting without model modifications. TIME-LLM (Jin et al., 2024) reprograms LLM by learning an alignment layer that translates time series values into its lexical space. Methods like LSTPrompt (Liu et al., 2024) prompt LLMs to reason separately about long and short range patterns. Despite these advances, their benefits over classical models remain uncertain. (Tang et al., 2024) shows that LLMs can handle periodic patterns but struggle with irregular ones, although external context can help. In user modeling, (Du et al., 2025)(Dong et al., ) incorporate irregular time intervals into LLMs for sequential recommendation, yet the task remains focused on what to recommend rather than when the event occurs.
@@ 2.1. Datasets
We evaluate our approach on two real-world e-commerce datasets: proprietary grocery data and the public Instacart dataset (M Yasser H., 2017). In both datasets, user-category pairs with fewer than five purchases are excluded, and intervals longer than 20 days are removed. The proprietary dataset contains 5,780 users across 12 product types (1.03 types/user average) with over 110,000 purchase records. The Instacart dataset includes 2,661 users across 10 categories (1.87 categories/user average) with roughly 194,000 orders. Both datasets span diverse categories (Fresh Produce, Dairy, Packaged Goods, Snacks) capturing variability in purchase periodicity and seasonality for temporal reasoning comparison.
@@ 3.1. Main Results
RQ1: LLMs vs. Traditional ML A primary finding from 1, consistent across both the proprietary and Instacart datasets, is the ML model’s dominant performance. It achieves over 50% improvement on most key metrics compared to the best-performing LLM. For instance, on the proprietary dataset, the ML model’s MAPE of 29.92% is 92.0% better than the best-performing LLM (Claude-3.5 Medium at 57.45%). On the business-critical ”TA@1” metric, the ML model scores 22.98%, beating the best LLM’s 14.63% (Claude-3.5 High) by 57.1%. This performance gap is likely due to a fundamental ”impedance mismatch” for the LLM. The ML model directly operates on structured numerical features, whereas the LLM must first translate qualitative linguistic descriptions into an internal representation before performing regression, introducing significant error. Notably, the disparity in performance is significantly larger for error metrics than for accuracy metrics, as LLMs’ performance on TA@k is competitive. This suggests that while LLMs fail at quantitative precision (i.e. pinpointing the exact day), they are better at approximate temporal classification. This aligns with the business goal for replenishment, where knowing a user will buy ”in the next 48 hours” is often sufficient.
@@
Another notable observation is that LLMs outperform the “stat best” median baseline across both datasets and metrics, indicating that they extract additional signal from contextual descriptions beyond simple historical central tendency. This suggests that LLMs incorporate contextual cues rather than merely reproducing medians, yielding more informative estimates, especially for approximation rather than exact-day prediction. These results indicate that LLMs have useful but limited temporal reasoning capabilities.
@@
RQ2: The Impact of Context on LLM Performance Our second research question examines how different levels of context affect LLM performance. Across both datasets, medium-level context, including basic product attributes, simple statistics, and concise user history, consistently improves both model metrics and business-oriented accuracy. In contrast, high-level narrative context often degrades performance, sometimes matching the zero-context baseline. This suggests a context-as-noise effect: focused, decision-relevant information aids temporal reasoning, whereas semantically rich but temporally irrelevant details distract the model. As illustrated in Figure 3, green highlights show that medium-context prompts encourage the model to rely on stable quantitative cues, producing correct or near-correct predictions, while red highlights show that high-context prompts surface narrative details that models overweight, diverting attention from the underlying temporal structure and leading to erroneous predictions.
[…]
@@
Our study presents a framework for time-interval prediction, establishes strong baselines, and positions LLMs between statistical and machine-learning models. We show that targeted and compact context can aid LLM reasoning, while excessive narrative detail can obscure the temporal signal. These findings highlight both the potential and limitations of LLMs for structured temporal inference and point toward hybrid models combining statistical precision with linguistic flexibility.
Figure 3. Examples of Claude-3.5-Sonnet prediction and reasoning with different levels of context.